OpenStudy (anonymous):

If only 1/4 of the amount of baking soda in the recipe were used how much baking soda would be used ? A. 1/64 B. 1/20 C. 1/16 D. 1/4 Help !!!

OpenStudy (anonymous):

is that the whole question?

OpenStudy (bibby):

there's obviously a recipe given but so many people just post partial questions and then leave

OpenStudy (anonymous):

@bibby yep.

OpenStudy (phi):

1/4 of the amount of baking soda here are some examples: 1/4 of 4 is 1 (notice 1/4 * 4 = 1) 1/4 of 8 is 2 (notice 1/4 * 8 = 8/4 = 2) 1/4 of "some number" means multiply that number by 1/4 or (the same thing) , divide the "some number" by 4 of course, to do your problem we need to know how much baking soda is used in the recipe